mathslover (mathslover):

If \(\tan (\pi \cos \theta) = \cot (\pi \sin \theta) \) , \(\theta \in \left( 0, \cfrac{\pi}{2} \right)\) , then \(\cos \left( \theta - \cfrac{\pi}{4} \right) = ? \) (Again, not a challenge)

mathslover (mathslover):

If \(\tan (\pi \cos \theta) = \cot (\pi \sin \theta) \) , \(\theta \in \left( 0, \cfrac{\pi}{2} \right)\) , then \(\cos \left( \theta - \cfrac{\pi}{4} \right) = ? \) (Again, not a challenge)

mathslover (mathslover):

What I like about this question is that we have tan and cot here, we can use them... or put sin and cos there instead of tan and cot.

mathslover (mathslover):

\(\cfrac{\sin (\pi \cos \theta) }{\cos (\pi \cos \theta) } = \cfrac{\cos (\pi \sin \theta)}{\sin (\pi \sin \theta) } \)

hartnn (hartnn):

\(\Large \tan(A+B) = ... have you come across this formula ?\)

OpenStudy (anonymous):

have you tried it yet kush?

mathslover (mathslover):

@mukushla yeah, these questions are in my notebook ... list of questions in which I need help. I tried it a lot. Tried to simplify, but I don't get anything.

mathslover (mathslover):

Do you want me to put that here? But, I really didn't get anything right.

hartnn (hartnn):

if tan A tan B = 1 A+B = pi/2

OpenStudy (goformit100):